OpenStudy (ofmiceparade):

Solve x2 − 8x + 5 = 0 using the completing-the-square method.

OpenStudy (ofmiceparade):

@hartnn Hi

hartnn (hartnn):

hello :) what have you tried?

OpenStudy (ofmiceparade):

i got x and square root of 11

OpenStudy (ofmiceparade):

I don't think i got it right

hartnn (hartnn):

ok, lets go step by step, x^2 -8x = -5 what did you add to both sides of this equation?

OpenStudy (ofmiceparade):

16

hartnn (hartnn):

good! so left side became the perfect square of (x-4) \((x-4)^2 = 16-5 \\ (x-4)^2=11\) now take square root on both sides

OpenStudy (ofmiceparade):

?

OpenStudy (ofmiceparade):

11 can't be square rooted

OpenStudy (ofmiceparade):

and idk about x-4

hartnn (hartnn):

right, keep it as \(\Large \sqrt{11}\)

OpenStudy (ofmiceparade):

okay

hartnn (hartnn):

\(\sqrt{(x-4)^2} = x-4\)

OpenStudy (ofmiceparade):

okay so x-4 = square root of 11

hartnn (hartnn):

\(\Large x-4 =\pm \sqrt {11}\)

hartnn (hartnn):

because there would be a negative root too

OpenStudy (ofmiceparade):

okay but my answer choices are x = negative four plus or minus the square root of eleven x = four plus or minus the square root of eleven x = negative four plus or minus the square root of five x = four plus or minus the square root of five

OpenStudy (ofmiceparade):

That isn't "x ="

hartnn (hartnn):

add 4 on both sides

OpenStudy (ofmiceparade):

OHHHH

hartnn (hartnn):

\(\Large x = 4\pm \sqrt{11}\)

OpenStudy (ofmiceparade):

OMG I LOVE YOU <3333333333333333333333333333333333

OpenStudy (ofmiceparade):

No hetero

hartnn (hartnn):

me too :) welcome ^_^